how can we develop the equtaion of potentila drop in the potential meter

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Potential drop is given as:
x = V/l
and also
V1/V2 = l1/l2
l would be the length measured where potentiometer reading shows zero.